# The Human Labour of Data Work: Capturing Cultural Diversity through World Wide Dishes

### v2

@@ 5.2. Data sources and reflective practice
    Documentation artefacts (created November 2023–April 2024): Throughout the development of the WWD data collection system, all COs, except the second author, contributed to documentation artefacts capturing the development and deployment of WWD. These include a living Google Doc where COs discussed community feedback, the onboarding Google Slide deck for CAs, as well as various documents for planning, collecting, and collating feedback. These documents captured real-time reflections during the development process.
@@
    Documentation artefacts (created November 2023–April 2024): Throughout the development of the WWD data collection system, all COs, except the second author, contributed to documentation artefacts capturing the development and deployment of WWD. These include a living Google Doc where COs discussed community feedback, the onboarding Google Slide deck for CAs, as well as various documents for planning, collecting, and collating feedback. These documents captured real-time reflections during the development process.
@@
    Formal paper write-up (May 2024): Around May 2024, the COs engaged in intentional reflection on the process of developing and deploying WWD, culminating in a paper submitted to a conference that focussed on the WWD dataset and its use for understanding representational harms in foundation models. During the paper writing process, the COs analysed the WWD database (e.g., regional coverage, number of contributions, interesting features of the data) and the documentation artefacts. As the COs continued to discuss the process of developing and deploying WWD between May and August, CAs and the role they played in making this project possible emerged as a key theme. The reflections and documents we produced during the formal paper write-up served as part of our reflective process for this CSCW paper. They helped us to crystallise our focus on CAs and their labour.
  4. (4)
    Post-mortem reflections (collected August–October 2024): Four months after data collection concluded, nine COs who also served as CAs produced structured reflections on their experiences. As we show in Figure 1, all COs served as CAs. Four COs provided audio recordings through voice notes in private WhatsApp conversations with the second author, who asynchronously asked follow-up questions similar to a semi-structured interview. These recordings were transcribed using MS Word’s voice-to-text feature. The remaining five COs wrote detailed reflections in a shared Google Doc. All reflections responded to prompts (see Appendix F for example of prompts) created by the second author, who has extensive experience with ethnographic research methods.
@@ 5.3. Analysis
Code Background Involvement in the African ML community Research Institution Community Ambassador Duration of involvement
@@
O6 Machine Learning Deep Learning Indaba, Deep Learning IndabaX Egypt Microsoft Egypt 8 months
@@ 5.4.2. How does who we are impact what we did?
Our positionality significantly shaped how we developed, deployed, reflected on, and made sense of the WWD project. As ML researchers engaging with CCs during the deployment process, we occupied positions of technical authority, which we leveraged to demonstrate how existing technologies inadequately represented local dishes, likely motivating community participation. Simultaneously, our identities as community members also likely played a significant role in establishing trust and buy-in. As community members, we had intimate knowledge of the relevant cultural norms which helped us to adapt our design of the WWD system to accommodate regional technological constraints and communication preferences.
Our use of first-person methods acknowledges the subjective nature of our experiences as CAs, with our analysis reflecting what we found significant—particularly the emotional and social labour involved in supporting WWD’s development. First-person methods allow us to recognise the produced and subjective nature of our reflections serving as CAs. The choice of things we focus on in our analysis is coloured by what we felt was important to unpack as community members: the things that weighed heavily on our consciences as we both orchestrated and supported the development and deployment of WWD. We embrace this subjectivity and treat it as an asset that helps us provide a rich, “thick description” (Geertz, 2008) of the social and emotional labor that participatory mediators perform in supporting participatory design research.
@@ 6. Building World Wide Dishes: Lessons from the field
Building the WWD dataset was a process that took over seven months, and involved twelve COs and more than 170 CAs and CCs. The process of launching the WWD data collection effort involved a beta-testing stage to solicit an initial round of feedback from community members about the data collection instrument and process, as well as hosting ongoing focus groups with CCs to surface issues with knowledge contributions. Through systematic analysis of post-mortems and field notes from the COs, we identify how CAs acted as participatory mediators who made this bottom-up, community-led data collection process possible. In this section, we identify three ways in which CAs made the WWD project possible by: building trust between community members and the research process, making participation accessible, and contextualising community values to support effective data collection.
@@ 6.1. Community Ambassadors build trust between community members and the research process
Collecting long-tailed data requires accessing and connecting with under-represented populations who tend to be “hard to reach”. [^12] We use quotation marks to draw attention to the fact that the term “hard to reach” is value-laden. Calling a population “hard to reach” redirects attention away from the reasons why that population may be difficult for the researcher to access and instead focuses attention on the group as the problem, rather than on the institution or researcher as the potential problem. For more, see (Benjamin, 2016; Epstein, 2008). Obtaining this access typically requires working with community gatekeepers who can serve as intermediaries between the researchers and community (Le Dantec and Fox, 2015; Epstein, 2008). However, the presence of a community gatekeeper [^13] We use the term “gatekeeper”, as it is known in participatory design literature, to refer to those community members who help outsiders interact with community members while preserving the integrity of the community (Epstein, 2008; Benjamin, 2016). in a project does not, in itself, resolve ethical questions around shifting the distribution of power from the researchers to the researched. In building WWD, we made an intentional choice to involve community gatekeepers (e.g., CAs) as members of the research team who shaped the data collection process, the development of the data collection tool, and the paper authorship process. This integration of CAs as core team members rather than mere gatekeepers established a foundation of trust, as community members could see their own representatives helping to shape the research rather than simply facilitating access to them.
The first author, who is a researcher at a prestigious Western university, recognised that their positionality put them in a position of power as relates to the collection of cultural data from communities to which they do not belong. While they are originally from South Africa and have been deeply involved in community efforts within the African ML community, specifically the Deep Learning Indaba, they felt that the WWD project must be co-led by members from communities who were contributing data. Therefore, they invited multiple people in their professional network who were active researchers and volunteers for local AI communities to be co-leads and co-authors on the project. Two of these, O5 and O7, are Nigerian and Sudanese respectively, and were experienced in conducting data collection efforts with communities within their home countries. As O5, O6, and O7 spread information about WWD to their networks, additional community members representing more than eight countries on the African continent signed up to be CAs for the project. CAs leveraged their positionality, relationships, and social capital to make data collection with communities in Africa possible.
@@
Established methods of accessing data contributors, such as cold-calling or simply distributing the call to participate on mass broadcast social media channels (e.g., X, formerly known as Twitter, LinkedIn), were arguably less successful, even when it was the CAs themselves who posted the call. Access to, or the ability to access, potential contributors was insufficient. Instead, CAs often had to rely on the relationships they had already established to solicit contributions. For example, O6 reported in their post-mortem that: “So, I kept sharing and sharing over social media … but the contributions were not growing as much as I hoped … So I reach out to people individually.” As a result, CAs tapped into their network of friends and family to solicit data contributions. However, reaching out to personal networks was not without risk.
@@
CAs controlled how they, and by proxy the larger WWD team, interacted with their communities. For example, the COs and CAs administered WhatsApp groups within the larger WWD WhatsApp community, where they carried out conversations with Contributors in the language of the region they oversaw. CAs had complete oversight over how they wanted to structure conversations within their WhatsApp group and enforce moderation rules that were proposed by the COs.
@@ 6.2. Community Ambassadors make participation accessible
The COs recognised that the CCs may face challenges in terms of Internet connectivity and would likely interact with our website via mobile devices, as also observed in the AYA community data collection initiative (Singh et al., 2024; Avle et al., 2018). Therefore, the COs devoted significant time to developing the website with specific considerations in mind to ensure the computational burden was not placed on the end-user device. Particular attention was paid to the responsiveness of the website to allow accessibility from different types of devices, and we can report that 62% of all CCs accessed the website from mobile devices as opposed to desktops. Additionally, in cases where Internet accessibility was not possible at all due to cellular network restrictions, as was the case for our CCs in the Democratic Republic of the Congo, the CA filled in the data on their behalf.
@@
Collecting image data for ML purposes, as WWD does, is a complex process due to concerns about image ownership. Because WWD was intended to be an open-source dataset with Creative Commons licensing, all submitted images either had to have a Creative Commons license (if it wasn’t an original) or be an original (e.g., taken by the contributor) photo shared with explicit permission that it could be used for research purposes. These guidelines were presented clearly on the data contribution form; however, CAs noticed early on that CCs were uploading images from the Internet that did not have Creative Commons licenses. As a result, the COs and CAs began hosting regular “teach-ins” during office hours for CCs to attend in order to learn how to properly make a data contribution. O6 explained the role of office hours in making data contribution guidelines accessible to participants:
“The [office hours] helped ensure our data collection met our requirements and that contributors understood both the licensing requirements and how to properly complete the submission forms, enabling us to collect exactly what we needed for the project” O6
@@
Office hours also served as a space for CCs to shape the research process. Office hours attendees were encouraged to share feedback about challenges they faced in making data contributions, which the CAs then brought to planning meetings with the COs. As a result, the COs amended the data collection instrument, produced additional guidelines for participation, and refined their recruitment strategy. O1 recalls how the team wanted to follow so-called “professional protocol” and communicate through official and professional media such as mailing lists and social media. The members of the community immediately asked for WhatsApp groups.
@@
“For the photos and information I couldn’t get myself, I asked certain resourceful people, particularly my parents and friends, to provide the photos and information about the utensils to use and the appropriate time to eat certain meals” O8, summarising what a CC told them
@@
“For the dishes I struggled to explain, I turned to various sources for help. When possible, I looked online to see how they were made. However, for some dishes, it was difficult to find accurate information. In those cases, I reached out to my mother and grandmother, who were able to share their expertise and guide me through the process” O8, summarising what a CC told them.
@@
Food is complex: dishes often harbour a deeper cultural significance. For example, O6 recalled how during their experience managing data contributions from Egypt, they realised that the classification of “Egyptian food” was fuzzy at best:
@@
For O6, capturing data about Egyptian cuisine led them to further investigate the origins of their own cultural dishes. In the weeks following the data collection effort, they attended a session about ”food as cultural currency” at the RiseUp Summit Egypt 2024 to learn more about the Egyptian cuisine through the eyes of local food experts.
@@
In post-mortem reflections, CAs highlighted the importance of recognising the limitations of using national borders to demarcate cultures. The granularity of representation was essential to ensuring that distinct cultures were not lumped together into a single flattened snapshot, as physical borders often do. O5 recalls how she realised that one of the three major ethnic groups in Nigeria—the Hausa community—was under-represented in the WWD dataset. They reached out to a personal connection whom they knew was a member of that ethnic community who, in turn, helped share the website with their community.
@@ 7. Discussion: Implications for Designing Infrastructure to Support Participatory AI Efforts
Drawing from our experience with WWD, particularly our interactions on the African continent, we surface five lessons for designing infrastructures that support a Participatory Turn in AI (Delgado et al., 2023). WWD is an example of how to build infrastructure for participatory AI that highlights the crucial role CAs play as participatory mediators. These mediators performed essential work connecting community members to the research initiative and from their efforts, we derive five key lessons for researchers building datasets for ML models using participatory approaches. These lessons guide how we might move towards more equitable structures of engagement with socially marginalised communities.
@@
Lesson Number One: Invest in and support participatory mediators. Collecting fine-grained and representative cultural data is exceptionally difficult. Cultures are not bounded by government borders and/or other manufactured, static systems, but rather may extend across regions and be the product of intercultural exchanges (Gupta and Ferguson, 2008). This, and the number of cultures present in any geographical area, makes knowing what cultural data is under-represented almost impossible without extensive consultation. As O1 detailed in their reflections, CAs play a crucial role in identifying what communities or regions are underserved. Their fine-grained understanding of community dynamics enables them to recognise who is systematically excluded or marginalised within these contexts. Then, CAs like O5 can tap into their insider identity to build trust and motivate participation, creating pathways for engagement with populations that might otherwise remain inaccessible to outside researchers. CAs therefore act as participatory mediators (Okamura et al., 1995) who play a crucial role not only in managing technical issues but also in understanding the contexts in which these data contribution efforts play out. Calls for equity-driven participatory design projects advocate researchers to consider the roles of historical context, research setting, and commitment to practical outcomes (Harrington et al., 2019; Reason and Bradbury, 2008). We address calls to actualise equity-driven participatory design engagements by developing the role of CAs who act as participatory mediators, contributing contextual knowledge about communities, bringing the research to communities, and using their insider status as community members to advocate for research goals in line with community interests. We call for future researchers to construct formal organisational roles within their projects for participatory mediators, ensuring their labour is both visible and properly supported within research infrastructures (Okamura et al., 1995; Grudin, 1988; Star and Strauss, 1999).
@@
Lesson Number Two: Researchers should build on existing community-led initiatives. Grassroots communities have a long history of engaging in data activism projects where community members utilise datafication as a tool to capture and legitimise their ways of knowing (e.g., epistemologies) (D’Ignazio, 2024; Milan and Velden, 2016; Collective, 1980; Harris, 2024). ML researchers seeking to mitigate breakdowns in model performance by collecting more data about socially marginalised groups should carefully investigate whether there are existing data activism efforts within communities. As D’Ignazio (2024) and Pierre et al. (2021) call for in their work, equity-driven participatory design engagements should centre the community by imagining the scope and purpose of data collection efforts—what do communities want to achieve in collecting data? Building tools for existing community-led initiatives is a step towards centring community ways of knowing in data activism efforts. In our work, we built upon existing research networks that aim to collect and record cultural data about socially marginalised groups, such as the Deep Learning Indaba and AI Saturdays Lagos (see Table 2 for a comprehensive list). Many of the authors of this paper, who served as CAs, are members of these research networks who see the value in (and potential pitfalls of) pursuing cultural preservation through digitisation. We advise that future researchers seeking to engage in participatory AI efforts, especially as they relate to developing datasets and ML systems for the African continent, familiarise themselves with the core values and missions of established local research networks and take steps to understand the space in which they exist and determine if there is room for non-extractive, mutually beneficial engagement. Doing so will enable researchers to better understand current priorities for community-led research and align their efforts to promote more equitable ML models with community definitions of equity and community criteria of success.
@@
Lesson Number Five: Examine the motivating factors that are presented to participants, and the promises made alongside them. Whilst better representation in datasets will improve quality of service to an extent, it is unlikely to undo the centuries of discrimination and identity prejudice that also influence these AI pipelines. Honest engagement can empower contributors to make informed decisions about how they choose to contribute their knowledge (Dalal et al., 2024). We intentionally chose not to make use of professional data centre workers [^14] Professional data centre workers are those people employed in a centralised manner to perform data collection tasks. Their livelihood is, therefore, connected to the requirement to engage in data contribution, which does not align with WWD goals. Additionally, even had we wished to use data centre workers, we lacked the resources to which a large technology company might have access, such as the ability to engage an ethical outsourcing company (e.g., Enlabler) to recruit and pay data workers., but instead used a data collection method that would enable us to reach participants other than those employed in a data worker centre, such as older generations and those across a wide socioeconomic range. Although we would have preferred to compensate each CC, we were unable to do so due to a lack of funding. We instead put out a call for volunteers, clearly stating the amount of work to be done, the motivation behind the project, and what we intended to do with the data after collection. CAs also demonstrated differences in the quality of service of widely available AI systems, such as under-representation of images and information returned by search queries, as means to encourage submissions. As a result, a number of individuals voluntarily signed up to contribute their time and effort to the project. However, despite recruitment successes, CAs also shared conflicting feelings about potentially exploitative and extractive processes when tapping into the intrinsic motivations that may have motivated CCs’ involvement. We surface this lesson because the majority of data contributions came from the African continent, and we asked why this might be.

### v3

@@ 5.2. Data sources and reflective practice
    Documentation artefacts (created November 2023–April 2024): Throughout the development of the WWD data collection system, all COs, except the second author, contributed to documentation artefacts capturing the development and deployment of WWD. These included a living Google Doc where COs discussed community feedback, the onboarding Google Slide deck for CAs, as well as various documents for planning, collecting, and collating feedback. These documents captured real-time reflections during the development process.
@@
    Formal paper write-up (May 2024): Around May 2024, the COs engaged in intentional reflection on the process of developing and deploying WWD, culminating in a paper submitted to a conference that focussed on the WWD dataset and its use for understanding representational harms in foundation models. During the paper writing process, the COs analysed the WWD database (e.g., regional coverage, number of contributions, interesting features of the data) and the documentation artefacts. As the COs continued to discuss the process of developing and deploying WWD, the CAs and the role they played in making this project possible emerged as a key theme. The reflections and documents we produced during the formal paper write-up served as part of our reflective process for this CSCW paper. They helped us to crystallise our focus on CAs and their labour.
@@
    Post-mortem reflections (collected August–October 2024): Four months after data collection concluded, nine COs who also served as CAs produced structured reflections on their experiences. As we show in Figure 1, all COs served as CAs. Four COs provided audio recordings through voice notes in private WhatsApp conversations with the second author, who asynchronously asked follow-up questions similar to a semi-structured interview. These recordings were transcribed using MS Word’s voice-to-text feature. The remaining five COs wrote detailed reflections in a shared Google Doc. All reflections responded to prompts (see Appendix F for full list of prompts) created by the second author, who has extensive experience with ethnographic research methods.
@@ 5.3. Analysis
Code Background Involvement in the African ML community Research Institution Community Ambassador Length of involvement
@@
O6 Machine Learning Deep Learning Indaba, Deep Learning IndabaX Egypt Independent Researcher Egypt 8 months
@@ 5.4.2. How does who we are impact what we did?
Our positionality significantly shaped how we developed, deployed, reflected on, and made sense of the WWD project. As ML researchers engaging with CCs during the deployment process, we occupied positions of technical authority, which we leveraged by demonstrating how existing technologies inadequately represented local dishes. This likely helped to motivate community participation. Simultaneously, our identities as community members also likely played a significant role in establishing trust and buy-in. As community members, we had intimate knowledge of the relevant cultural norms which helped us to adapt our design of the WWD system to accommodate regional technological constraints and communication preferences.
Our use of first-person methods acknowledges the subjective nature of our experiences as CAs, with our analysis reflecting what we found significant—particularly the emotional and social labour involved in supporting WWD’s development. First-person methods allow us to recognise the produced and subjective nature of our reflections serving as CAs. The choice of topics we focus on in our analysis is coloured by what we felt was important to unpack as community members: the things that weighed heavily on our consciences as we both orchestrated and supported the development and deployment of WWD. We embrace this subjectivity and treat it as an asset that helps us provide a rich, “thick description” (Geertz, 2008) of the social and emotional labor that participatory mediators perform in supporting participatory design research.
@@ 6. Building World Wide Dishes: Lessons from the field
Building the WWD dataset was a process that took over seven months, and involved twelve COs and more than 170 CAs and CCs. The process of launching the WWD data collection effort involved a beta-testing stage to solicit an initial round of feedback from community members about the data collection instrument and process, as well as hosting ongoing focus groups with CCs to surface issues with knowledge contributions. Through systematic analysis of post-mortems and field notes from the COs, we identify how CAs acted as participatory mediators who made this bottom-up, community-led data collection process possible. In this section, we identify three ways in which CAs made the WWD project possible: building trust between community members and the research process, making participation accessible, and contextualising community values to support effective data collection.
@@ 6.1. Community Ambassadors build trust between community members and the research process
Collecting long-tailed data requires accessing and connecting with underrepresented populations who tend to be “hard to reach”. [^11] We use quotation marks to draw attention to the fact that the term “hard to reach” is value-laden. Calling a population “hard to reach” redirects attention away from the reasons why that population may be difficult for the researcher to access and instead focuses attention on the group as the problem, rather than on the institution or researcher as the potential problem. For more, see (Benjamin, 2016; Epstein, 2008). Obtaining this access typically requires working with community gatekeepers who can serve as intermediaries between the researchers and community (Le Dantec and Fox, 2015; Epstein, 2008). However, the presence of a community gatekeeper [^12] We use the term “gatekeeper”, as it is known in participatory design literature, to refer to those community members who help outsiders interact with community members while preserving the integrity of the community (Epstein, 2008; Benjamin, 2016). in a project does not, in itself, resolve ethical questions around shifting the distribution of power from the researchers to the researched. In building WWD, we made an intentional choice to involve community gatekeepers (e.g., CAs) as members of the research team who shaped the data collection process, the development of the data collection tool, and the paper authorship process. This integration of CAs as core team members rather than mere gatekeepers established a foundation of trust, as community members could see their own representatives helping to shape the research rather than simply facilitating access to them.
The first author, who is a researcher at a prestigious Western university, recognised that their positionality put them in a position of power as relates to the collection of cultural data from communities to which they do not belong. While they are originally from South Africa and have been deeply involved in community efforts within the African ML community, they felt that the WWD project must be co-led by members from communities who were contributing data. Therefore, they invited multiple people in their professional network who were active researchers and volunteers for local AI communities to be co-leads and co-authors on the project. Two of these, O5 and O7, are Nigerian and Sudanese respectively, and were experienced in conducting data collection efforts with communities within their home countries. As O5, O6, and O7 spread information about WWD to their networks, additional community members representing more than eight countries on the African continent signed up to be CAs for the project. CAs leveraged their positionality, relationships, and social capital to make data collection with communities in Africa possible.
@@
Established methods of accessing data contributors, such as cold-calling or simply distributing the call to participate on mass broadcast social media channels (e.g., X, formerly known as Twitter, LinkedIn), were arguably less successful, even when it was the CAs themselves who posted the call. Access to, or the ability to access, potential contributors was insufficient. Instead, CAs often had to rely on the relationships they had already established to solicit contributions. For example, O6 reported in their post-mortem that: “So, I kept sharing and sharing over social media … but the contributions were not growing as much as I hoped … So I reach out to people individually.” As a result, CAs tapped into their network of friends and family to solicit data contributions. However, reaching out to personal networks was not without risk, as CAs were making potentially uncomfortable requests for volunteer work from friends and relatives.
@@
CAs controlled how they, and by proxy the larger WWD team, interacted with their communities. For example, the COs and CAs administered WhatsApp groups within the larger WWD WhatsApp community, where they carried out conversations with CCs in the language of the region they oversaw. CAs had complete oversight over how they wanted to structure conversations within their WhatsApp group and enforce moderation rules that were proposed by the COs.
@@ 6.2. Community Ambassadors make participation accessible
The COs recognised that the CCs may face challenges in terms of Internet connectivity and would likely interact with our website via mobile devices, as also observed in the AYA community data collection initiative (Singh et al., 2024; Avle et al., 2018). Therefore, the COs devoted significant time to developing the website with specific considerations in mind to ensure the computational burden was not placed on the end-user device. Particular attention was paid to the responsiveness of the website to allow accessibility from different types of devices, and we can report that 62% of all CCs accessed the website from mobile devices as opposed to desktops. Additionally, in cases where Internet accessibility was not possible at all due to cellular network restrictions, as was the case for our CCs in the Democratic Republic of the Congo, the CA filled in their data contributions on their behalf.
@@
Collecting image data for ML purposes, as WWD does, is a complex process due to concerns about image ownership. Because WWD was intended to be an open-source dataset with Creative Commons licensing, all submitted images either had to have a Creative Commons license (if it wasn’t an original) or be an original (e.g., taken by the contributor) photo shared with explicit permission that it could be used for research purposes. These guidelines were presented clearly on the data contribution form; however, CAs noticed early on that CCs were uploading images from the Internet that did not have Creative Commons licenses. As a result, the COs and CAs began hosting regular “teach-ins” during “office hours” for CCs to attend in order to learn how to properly make a data contribution. O6 explained the role of office hours in making data contribution guidelines accessible to participants:
“The [office hours] helped ensure our data collection met our requirements and that contributors understood both the licensing requirements and how to properly complete the submission forms, enabling us to collect exactly what we needed for the project.” O6
@@
Office hours also served as a space for CCs to shape the research process. Office hours attendees were encouraged to share feedback about challenges they faced in making data contributions, which the CAs then fed back to the COs. As a result, the COs amended the data collection instrument, produced additional guidelines for participation, and refined their recruitment strategy. O1 recalls how the team wanted to follow so-called “professional protocol” and communicate through official and professional media such as mailing lists and social media. The members of the community immediately asked for WhatsApp groups.
@@
“For the photos and information I couldn’t get myself, I asked certain resourceful people, particularly my parents and friends, to provide the photos and information about the utensils to use and the appropriate time to eat certain meals.” O8, summarising what a CC told them
@@
“For the dishes I struggled to explain, I turned to various sources for help. When possible, I looked online to see how they were made. However, for some dishes, it was difficult to find accurate information. In those cases, I reached out to my mother and grandmother, who were able to share their expertise and guide me through the process.” O8, summarising what a CC told them
@@
Food is complex and dishes often harbour a deeper cultural significance. For example, O6 recalled how during their experience managing data contributions from Egypt, they realised that the classification of “Egyptian food” was fuzzy at best:
@@
For O6, capturing data about Egyptian cuisine led them to further investigate the origins of their own cultural dishes. In the weeks following the data collection effort, they attended a session about ”food as cultural currency” at the RiseUp Summit Egypt 2024 to learn more about Egyptian cuisine through the eyes of local food experts.
@@
In post-mortem reflections, CAs highlighted the importance of recognising the limitations of using national borders to demarcate cultures. The granularity of representation was essential to ensuring that distinct cultures were not lumped together into a single flattened snapshot, as physical borders often do. O5 recalls how she realised that one of the three major ethnic groups in Nigeria—the Hausa community—was underrepresented in the WWD dataset. They reached out to a personal connection whom they knew was a member of that ethnic community who, in turn, helped share the website with their community.
@@ 7. Discussion: Implications for Designing Infrastructure to Support Participatory AI Efforts
Drawing from our experience with WWD, particularly our interactions on the African continent, we surface five lessons for designing infrastructures that support a Participatory Turn in AI. WWD is an example of how to build infrastructure for participatory AI that highlights the crucial role CAs play as participatory mediators. These mediators performed essential work connecting community members to the research initiative and from their efforts, we derive five key lessons for researchers building datasets for ML models using participatory approaches. These lessons guide how we might move towards more equitable structures of engagement with socially marginalised communities.
@@
Lesson Number One: Invest in and support participatory mediators. Collecting fine-grained and representative cultural data is exceptionally difficult. Cultures are not bounded by government borders and/or other manufactured, static systems, but rather may extend across regions and be the product of intercultural exchanges (Gupta and Ferguson, 2008). This, and the number of cultures present in any geographical area, makes knowing what cultural data is underrepresented almost impossible without extensive consultation. As O1 detailed in their reflections, CAs play a crucial role in identifying what communities or regions are underserved. Their fine-grained understanding of community dynamics enables them to recognise who is systematically excluded or marginalised within these contexts. Then, CAs like O5 can tap into their insider identity to build trust and motivate participation, creating pathways for engagement with populations that might otherwise remain inaccessible to outside researchers. CAs therefore act as participatory mediators (Okamura et al., 1995) who play a crucial role not only in managing technical issues but also in understanding the contexts in which these data contribution efforts play out. Calls for equity-driven participatory design projects advocate researchers to consider the roles of historical context, research setting, and commitment to practical outcomes (Harrington et al., 2019; Reason and Bradbury, 2008). We address calls to actualise equity-driven participatory design engagements by developing the role of CAs who act as participatory mediators, contributing contextual knowledge about communities, bringing the research to communities, and using their insider status as community members to advocate for research goals in line with community interests. We call for future researchers to construct formal organisational roles within their projects for participatory mediators, ensuring their labour is both visible and properly supported within research infrastructures (Okamura et al., 1995; Grudin, 1988; Star and Strauss, 1999).
@@
Lesson Number Two: Researchers should build on existing community-led initiatives. Grassroots communities have a long history of engaging in data activism projects where community members utilise datafication as a tool to capture and legitimise their ways of knowing (e.g., epistemologies) (D’Ignazio, 2024; mil, [n. d.]; Collective, 1980; Harris, 2024). ML researchers seeking to mitigate breakdowns in model performance by collecting more data about socially marginalised groups should carefully investigate whether there are existing data activism efforts within communities. As D’Ignazio (2024) and Pierre et al. (2021) call for in their work, equity-driven participatory design engagements should centre the community by imagining the scope and purpose of data collection efforts—what do communities want to achieve in collecting data? Building tools for existing community-led initiatives is a step towards centring community ways of knowing in data activism efforts. In our work, we built upon existing research networks that aim to collect and record cultural data about socially marginalised groups, such as the Deep Learning Indaba and AI Saturdays Lagos (see Table 2 for a comprehensive list). Many of the authors of this paper, who served as CAs, are members of these research networks who see the value in (and potential pitfalls of) pursuing cultural preservation through digitisation. We advise that future researchers seeking to engage in participatory AI efforts, especially as they relate to developing datasets and ML systems for the African continent, familiarise themselves with the core values and missions of established local research networks and take steps to understand the space in which they exist and determine if there is room for non-extractive, mutually beneficial engagement. Doing so will enable researchers to better understand current priorities for community-led research and align their efforts to promote more equitable ML models with community definitions of equity and community criteria of success.
@@
Lesson Number Five: Examine the motivating factors that are presented to participants, and the promises made alongside them. Whilst better representation in datasets will improve quality of service to an extent, it is unlikely to undo the centuries of discrimination and identity prejudice that also influence these AI pipelines. Honest engagement can empower contributors to make informed decisions about how they choose to contribute their knowledge (Dalal et al., 2024). We intentionally chose not to make use of professional data centre workers, [^13] Professional data centre workers are those people employed in a centralised manner to perform data collection tasks. Their livelihood is, therefore, connected to the requirement to engage in data contribution, which does not align with WWD goals. Additionally, even had we wished to use data centre workers, we lacked the resources to which a large technology company might have access, such as the ability to engage an ethical outsourcing company (e.g., Enlabler) to recruit and pay data workers. but instead used a data collection method that would enable us to reach participants other than those employed in a data worker centre, such as older generations and those across a wide socioeconomic range. Although we would have preferred to compensate each CC, we were unable to do so due to a lack of funding. We instead put out a call for volunteers, clearly stating the amount of work to be done, the motivation behind the project, and what we intended to do with the data after collection. CAs also demonstrated differences in the quality of service provided by widely available AI systems, such as underrepresentation of images and information returned by search queries, as means to encourage submissions. As a result, a number of individuals voluntarily signed up to contribute their time and effort to the project. However, despite recruitment successes, CAs also shared conflicting feelings about potentially exploitative and extractive processes when tapping into the intrinsic motivations that may have motivated CCs’ involvement. We surface this lesson because the majority of data contributions came from the African continent, and we asked why this might be.